# Using Causality for Enhanced Prediction of Web Traffic Time Series

### v1

@@ Abstract.
Predicting web service traffic has significant social value, as it can be applied to various practical scenarios, including but not limited to dynamic resource scaling, load balancing, system anomaly detection, service-level agreement compliance, and fraud detection. Web service traffic is characterized by frequent and drastic fluctuations over time and are influenced by heterogeneous web user behaviors, making accurate prediction a challenging task. Previous research has extensively explored statistical approaches, and neural networks to mine features from preceding service traffic time series for prediction. However, these methods have largely overlooked the causal relationships between services. Drawing inspiration from causality in ecological systems, we empirically recognize the causal relationships between web services. To leverage these relationships for improved web service traffic prediction, we propose an effective neural network module, CCMPlus, designed to extract causal relationship features across services. This module can be seamlessly integrated with existing time series models to consistently enhance the performance of web service traffic predictions. We theoretically justify that the causal correlation matrix generated by the CCMPlus module captures causal relationships among services. Empirical results on real-world datasets from Microsoft Azure, Alibaba Group, and Ant Group confirm that our method surpasses state-of-the-art approaches in Mean Squared Error (MSE) and Mean Absolute Error (MAE) for predicting service traffic time series. These findings highlight the efficacy of leveraging causal relationships for improved predictions.
Figure 1. This figure illustrates 7-day search interest data for Netflix and Microsoft Outlook from Google Trends. The vertical axis represents normalized search interest as a percentage relative to the highest point on the chart for the specified time period. The horizontal axis corresponds to the time period from January 5th, 2025 to January 12th, 2025.
Figure 2. Illustration of ecological and analogical web service causality. The rabbit population influences grass abundance [19], where an increase in rabbits leads to a decrease in grass. This principle inspires our CCMPlus module, which captures causal effects among web services.
1. Introduction
@@
User-oriented web services continue to grow exponentially, especially with the advancements in artificial intelligence (AI) techniques [37; 26; 41], which have significantly accelerated the development of a diverse range of customized web applications. These services attract a substantial user base and play a pivotal role in enabling various social and practical activities. For instance, YouTube has amassed 2.7 billion users  [^1] https://www.globalmediainsight.com/blog/youtube-users-statistics/, powered by its cutting-edge recommendation algorithms. Accurately predicting web service traffic carries significant social and practical value, with applications spanning dynamic resource scaling [30; 53], load balancing [31], system anomaly detection [27], service-level agreement compliance [24], fraud detection [3], and so on. These capabilities not only enhance system performance but also improve the overall user experience with these technologies [23].
@@
Co-located, long-running web services often experience diverse workload patterns [53]. Web service traffic are characterized by frequent and significant fluctuations over time, driven by heterogeneous user behaviors at any given moment. These factors collectively make predicting web service traffic a highly challenging task [38; 30]. Previous works conduct extensive research to predict the web service traffic, often formulating it as a typical time series forecasting task. These approaches can broadly be categorized into statistical [21; 15; 22], machine learning [18; 7; 16], and deep learning [35; 12; 30; 53] methods. While statistical methods struggle to handle multi-dimensional and non-linear traffic data, machine learning methods address these limitations but fail to achieve the same level of accuracy as deep learning approaches. Among the deep learning based methods, the recent advancements in Transformer [43] architecture have demonstrated superior performance in sequential prediction tasks. Consequently, Transformer-based methods have emerged, achieving promising results [33] in web service traffic prediction. However, they did not pay enough attention to the causal relationship across web services.
In addition to the three types of specialized methods previously discussed for predicting web service traffic, general time series forecasting methods are also widely employed in practice [53; 4]. These approaches are typically divided into two categories [34]: statistical methods and neural network-based methods. Since statistical methods often fail to capture complex temporal features, neural network-based methods generally achieve superior performance [48]. Neural network-based approaches can be further classified into five paradigms [45; 40]: Recurrent neural network (RNN)-based, convolutional neural network (CNN)-based, Transformer-based, multi-layer perceptron (MLP)-based, and large language model (LLM)-based methods. CNN-based methods [44; 14] use convolutional kernels along the temporal dimension to identify sequential patterns, whereas RNN-based methods [11; 10; 32] rely on recurrent structures to model temporal state transitions. Transformer-based methods [52; 25; 6; 47] are widely recognized for their ability to effectively extract features using attention mechanisms. LLM-based methods [40; 42; 20] leverage advanced reasoning capabilities by processing time series data through specially designed prompts, offering a promising avenue for temporal modeling. Finally, MLP-based methods [47; 28; 8; 48] strike a compelling balance between forecasting accuracy and computational efficiency. It is regrettable that when advanced general time series forecasting methods are applied to web traffic prediction, the causal relationships among services are not utilized.
@@
Drawing inspiration from ecological causality, as illustrated in Figure 2, where grass abundance and rabbit populations influence one another iteratively [19], we identify analogous patterns in web service traffic as shown in Figure 1. Empirical observations of Google Trends data reveal a causal relationship between leisure websites, such as Netflix, and work-related software, like Outlook. Specifically, increased web traffic to Netflix corresponds to decreased traffic to Outlook, and vice versa. These empirical observations suggest the existence of latent causal relationships underlying human-driven web behaviors. By uncovering and leveraging these causal relationships, we can achieve more accurate modeling of service traffic patterns. Building on this motivation and the causality theory of Convergent Cross Mapping (CCM) [39], which originates from ecology, we introduce the CCMPlus module. This module extracts features from web service traffic time series while including causal relationships among services, thereby enhancing the accuracy of web service traffic prediction. Furthermore, the CCMPlus module could integrate easily with existing time series forecasting models, enriching them with more informative, causally-aware features.
Concretely, we extend the CCM theory to effectively merge with neural networks, resulting in the development of the CCMPlus module. The CCMPlus module operates in three key steps: first, it extracts initial feature representations from web service traffic time series; second, it computes a causal correlation matrix from a multi-manifold space based on these feature representations; and finally, it applies the causal correlation matrix to the initial feature representations, generating a resulting feature representation that incorporates informative causal information. This enhanced feature representation can then be concatenated with the feature representations of web service traffic time series extracted by existing time series models, thereby improving prediction accuracy.
Our main contributions in this work can be summarized as follows:
@@
  • Method: The CCMPlus module enhances existing time series forecasting models by generating feature representations that incorporate causal relationships across web services, addressing a critical limitation of many previous methods. Additionally, the CCMPlus module is designed for seamless integration with existing time series forecasting models, further contributing to improved prediction accuracy.
@@
  • Experiments: Experiments conducted on three real-world web service traffic datasets (Alibaba Group, Microsoft Azure, and Ant Group) demonstrate that our method achieves superior performance in terms of Mean Squared Error (MSE) and Mean Absolute Error (MAE) compared to previous state-of-the-art (SOTA) methods, thereby validating the effectiveness of the CCMPlus module.
The remainder of the paper is organized as follows. Section 2 reviews related work. Section 3 provides preliminary information about our proposed method. Section 4 elaborates on the proposed framework, encompassing the CCMPlus and time series backbone modules. Section 5 compares experimental results from the proposed method and prevalent methods, and analyzes the method and results. Section 6 concludes this paper.
2. Related Work
2.1. Web Service Traffic Prediction
@@
Web service traffic prediction is a task critical to enabling service autoscaling, load balancing, and anomaly detection. As web service traffic is often represented as time series data, existing approaches primarily frame traffic prediction as a time series prediction problem [30; 53; 4].
@@
Early research focused on statistical prediction methods [21; 15; 22] such as Moving Average, Auto-Regression, and Autoregressive Integrated Moving Average. These methods are valued for their simplicity and interpretability but are constrained by strict stationarity requirements. Moreover, they struggle to extend to multi-dimensional or non-linear data, limiting their applicability in dynamic environments. To overcome these limitations, methods like HOPBLR [18], LLR [7] and TWRES [16] employed machine learning techniques, including Logistic Regression and Support Vector Regression, respectively, for traffic prediction. However, these approaches were hindered by the limited expressive capacity of their models, resulting in suboptimal prediction accuracy. More recent advancements have shifted towards deep learning methods. CrystalLP [35] and GRUWP [12] utilize Long Short-Term Memory networks and Gated Recurrent Units, respectively, to predict service workloads. MagicScaler [30] proposes a novel multi-scale attentive Gaussian process-based predictor, capable of accurately forecasting future demands by capturing scale-sensitive temporal dependencies. The Performer [33] integrates the Transformer architecture into an encoder-decoder paradigm for service workload prediction. It leverages the self-attention mechanism to model temporal correlations and learn both global and local representations effectively. OptScaler [53] advances this direction with a proactive prediction module comprising a long-term periodic block and a short-term local block to capture multi-scale temporal dependencies. While existing traffic prediction methods demonstrate high accuracy through advanced time series forecasting techniques, they largely neglect the underlying causal relationships within the web services. Exploring these hidden causalities holds significant potential for further improving prediction performance.
2.2. Time Series Forecasting
@@
Besides approaches specifically tailored for web service traffic prediction, general time series forecasting methods are also applied to predict web traffic [4; 53; 9].
@@
Recent advancements in neural network architectures have significantly enhanced temporal modeling capabilities. Neural network approaches for time series forecasting can be categorized into five paradigms [45; 40]: RNN-based, CNN-based, Transformer-based, MLP-based, and large language model (LLM)-based methods.
Empirical methods often integrate components from the aforementioned categories, utilizing specific designs to effectively capture critical temporal features [45]. These specific designs incorporate series decomposition, multi-periodicity analysis, and multi-scale mixing architectures.
@@
For series decomposition, Autoformer [49] introduces a decomposition block based on moving averages, enabling the separation of complex temporal variations into seasonal and trend components. Building on this foundation, DLinear [50] utilizes series decomposition as a preprocessing step prior to performing linear regression. Crossformer [51] segments time series data into subseries-level patches and employs a Two-Stage Attention layer to effectively model cross-time and cross-variable dependencies within each patch. iTransformer [25] leverages the global representation of entire series and applies attention mechanisms to these series-wise representations, facilitating the capture of multivariate correlations. TimeXer [47] integrates external information into the Transformer architecture through a carefully designed embedding strategy, allowing the inclusion of external information into patch-wise representations of endogenous series. In the context of multi-periodicity, NBEATS [29] employs multiple trigonometric basis functions to model time series, providing a robust framework for handling periodic patterns. Similarly, TimesNet [48] applies Fourier Transform to decompose time series into components of varying periodic lengths and utilizes a modular architecture to process these decomposed components effectively. With respect to multi-scale mixing architectures, Pathformer [6] adopts multi-scale patch representations and applies dual attention mechanisms across these patches to capture both global correlations and local details, thereby addressing temporal dependencies comprehensively. TimeMixer [45] captures temporal features by introducing a novel multi-scale mixing architecture, which comprises two key components: Past-Decomposable-Mixing, designed to leverage disentangled series for multi-scale representation learning, and Future-Multipredictor-Mixing, which ensembles complementary forecasting skills across multi-scale series to enhance prediction accuracy.
@@
While general time series forecasting methods have been applied to web service traffic prediction [30; 53; 5], these methods often overlook causal relationships between services. In contrast, our CCMPlus module computes a causal correlation matrix used to generate temporal features incorporating causal relationships, significantly improving web service traffic prediction accuracy.
Figure 3. Web service traffic time series from the Microsoft Azure cluster. The horizontal axis represents time progression, while the vertical axis indicates the number of service requests at each time point.
3. Preliminaries
@@
In this section, we begin by presenting the time series patterns observed in web service traffic, illustrated with a specific web service traffic example. Subsequently, we provide an overview of the ecological causality theory, Convergent Cross Mapping (CCM) [39], which serves as the theoretical foundation for our proposed CCMPlus module for web service traffic prediction.
3.1. Web Service Traffic
@@
Figure 3 presents an illustrative example of web service traffic time series from Microsoft Azure. It exhibits significant fluctuations and frequent changes, primarily driven by human behavior and activity patterns. The web traffic prediction task can be formulated as follows:
$y(t)=P(y(t-\alpha),y(t-2\alpha),\dots,y(t-k\alpha)),$
where $y(t)$ is the number of request at time $t$, $\alpha$ denotes the prediction granularity, and $k$ is the historicial sequence length. However, accurately predicting web service traffic remains a well-recognized challenge due to the inherent complexity of traffic patterns, as highlighted by the research community [30; 53].
3.2. Convergent Cross Mapping
@@
Convergent Cross Mapping (CCM) theory published in Science [39] was originally proposed in the field of ecology and is designed to detect causal relationships between species. The exposition of CCM is often illustrated using the context of the Lorenz system, as depicted in Figure 4.
As depicted in Figure 4, the trajectory of the Lorenz system forms a manifold $M$ in the state space. This manifold $M$ consists of a collection of points that represent all possible states of the Lorenz system over time, with these points connected to create a structured geometric space. The manifold, also referred to as the attractor, encompasses all trajectories and potential states $\underline{m}(t)$ of the system. Each state $\underline{m}(t)$ corresponds to a point in $M$, represented by the coordinate vector $\underline{m}(t)=[X(t),Y(t),Z(t)]$.
@@
The shadow manifold, $M_{x}$ or $M_{y}$, represents the projection of the original manifold $M$ onto the system variables $X$ or $Y$, respectively. Specifically, a lagged coordinate embedding utilizes $E$ time-lagged values of $X(t)$ as coordinate axes to reconstruct the shadow manifold $M_{x}$. A point on $M_{x}$, denoted as $\underline{x}(t)$, is an $E$-dimensional vector expressed as:
@@
where $\tau$ is a positive time lag, and $E$ denotes the embedding dimension. In Figure 4, $E=3$. Similarly, the same approach applies to points $\underline{y}(t)$ in the manifold $M_{y}$, defined as:
$\underline{y}(t)=[Y(t),Y(t-\tau),Y(t-2\tau),\dots,Y(t-(E-1)\tau)].$
3.2.1. Cross Mapping
Cross mapping refers to the process of identifying contemporaneous points in the manifold $M_{x}$ of one variable $X$ based on points in the manifold $M_{y}$ of another variable $Y$. Specifically, given a point $\underline{y}(t)$ in the manifold $M_{y}$, the corresponding point in time from the manifold $M_{x}$ is $\underline{x}(t)$.
@@ 3.2.1. Cross Mapping
As illustrated in Figure 5, if $Y$ exerts a causal effect on $X$, information from $Y$ will be stored in $X$. Consequently, the neighbors of $\underline{x}(t)$ in $M_{x}$ will correspond to points with the same time indices in $M_{y}$, and these corresponding points will also be neighbors of $\underline{y}(t)$. However, as noted in [39], if $Y$ has no causal effect on $X$, the information about $Y$ in $X$ will be incomplete. As a result, the timely corresponding points in $M_{y}$ will diverge and no longer be neighbors of $\underline{y}(t)$.
Figure 4. The Convergent Cross Mapping (CCM) theory is often explained using the Lorenz system, utilizing its original manifold and corresponding shadow manifolds.
Figure 5. Cross mapping. The point $\underline{y}(t)$ in the manifold $M_{y}$ corresponds to the contemporaneous point in time $\underline{x}(t)$ in the manifold $M_{x}$.
Figure 6. Convergent predictability as the time series length increases, assuming $Y$ has a causal effect on $X$.
3.2.2. Convergence
@@
Convergence in CCM implies that if the variable $Y$ has a causal effect on $X$, extending the observation period improves the ability to predict $Y$ using points on the shadow manifold $M_{x}$, as illustrated in Figure 6. A longer observation period provides more trajectories to fill the gaps in the manifold, resulting in a more defined structure, which enhances the prediction of $\hat{Y}(t)\mid M_{x}$. Conversely, if two variables do not have a causal relationship, refining their manifolds will not lead to an improvement in predictive accuracy.
3.2.3. CCM Procedure
    $\hat{Y}(t)\mid M_{X}=\sum_{i=1}^{E+1}w_{i}Y(t_{i}),$
[…]
    $w_{i}=\frac{u_{i}}{\sum_{j=1}^{E+1}u_{j}},$
    where
4. Methodology

### v2

# Leveraging Latent Causal Relationships Among Web Services for Traffic Prediction

@@ Abstract
Predicting web service traffic is crucial for system operation tasks including dynamic resource scaling, anomaly detection, and fraud detection. Web service traffic is characterized by frequent and drastic fluctuations over time and are influenced by heterogeneous user behaviors, making accurate prediction a challenging task. Previous research has extensively explored statistical approaches, and neural networks to mine features from preceding service traffic time series for prediction. However, these methods have largely overlooked the latent causal relationships between services. Drawing inspiration from causality in ecological systems, we empirically recognize the causal relationships between web services. To leverage these relationships for improved traffic prediction, we propose an effective neural network module, CCMPlus, designed to extract causal relationship features across services. This module can be seamlessly integrated with existing time series models to consistently enhance the performance of traffic predictions. We theoretically justify that the causal correlation matrix generated by the CCMPlus module captures causal relationships among services. Empirical results on real-world datasets from Microsoft Azure, Alibaba Group, and Ant Group confirm that our method surpasses state-of-the-art approaches in Mean Squared Error and Mean Absolute Error for predicting service traffic time series. These findings highlight the efficacy of feature representations from the CCMPlus module.
[…]
6754
1 Introduction
@@
User-oriented web services continue to grow exponentially, which has significantly accelerated the development of a diverse range of customized web applications [20, 34, 36, 35, 37]. These services attract a substantial user base and play a pivotal role in enabling various social and practical activities. For instance, YouTube has amassed 2.7 billion users [^2] https://www.globalmediainsight.com/blog/youtube-users-statistics/, powered by its cutting-edge recommendation algorithms. Accurately predicting web service traffic carries significant social and practical value, with applications spanning dynamic resource scaling [25, 48], load balancing [26], anomaly detection [23], service-level agreement compliance [21], fraud detection [2], and so on. These capabilities not only enhance system performance but also improve the overall user experience [19].
@@
Co-located, long-running web services often experience diverse workload patterns [48]. Web service traffic is characterized by frequent and significant fluctuations over time, driven by heterogeneous user behaviors. These factors collectively make predicting web service traffic a highly challenging task [31, 25]. Previous works formulates this task as a typical time series forecasting task. These approaches can broadly be categorized into statistical [17, 11, 18], machine learning [14, 7, 12], and deep learning [29, 9, 25, 48] methods. While statistical methods struggle to handle multi-dimensional and non-linear traffic data, machine learning methods address these limitations but fail to achieve the same level of accuracy as deep learning approaches [48, 3]. The recent advancements in Transformer [39] architecture have demonstrated superior performance in sequential prediction tasks. Consequently, Transformer-based methods [27] have emerged, achieving promising results in web service traffic prediction. Aside from these ad-hoc models, general state-of-the-art (SOTA) Transformer-based time-series prediction methods [47, 22, 6, 42] also have the potential to be applied to the web traffic task. Recently emerged LLM-based methods [33, 38, 16] leverage advanced reasoning capabilities by processing time series data through specially designed prompts, offering a promising avenue for temporal modeling.
Figure 1: This figure illustrates the 7-day normalized search interest data for Netflix and Microsoft Outlook from Google Trends in the UK. The two services exhibit a causal relationship: an increase in Netflix usage corresponds to a decrease in the use of the email service.
@@
Drawing inspiration from ecological causality, for example, grass abundance and rabbit populations influencing one another iteratively [15], we identify analogous patterns in web service traffic as shown in Figure 1. Empirical observations of Google Trends data reveal a causal relationship between leisure websites, such as Netflix, and work-related software, like Outlook. Specifically, increased web traffic to Netflix corresponds to decreased traffic to Outlook, and vice versa. These empirical observations suggest the existence of latent causal relationships underlying human-driven web behaviors. By uncovering and leveraging these causal relationships, we can achieve more accurate modeling of service traffic patterns. Building on this motivation and the causality theory of Convergent Cross Mapping (CCM) [32], which originates from ecology, we introduce the CCMPlus module. This module extracts features from web service traffic time series while including causal relationships among services, thereby enhancing the accuracy of web service traffic prediction. Furthermore, the CCMPlus module could integrate easily with existing time series forecasting models, enriching them with more informative, causally-aware features.
Concretely, we extend the CCM theory to effectively merge with neural networks, resulting in the development of the CCMPlus module. Specifically, the traditional CCM theory relies on a set of expert designed hyper-parameters, which might introduce severe bias to the causality identification. To alleviate this, we propose a multi-manifold embedding space which is constructed by learnable parameters and present the causality relationship in diverse spaces. After obtaining the intial embedding, we employ the CCM procedure but in the multi-manifold space to calculate the causal correlation matrix, which is updated with a momentum update mechanism. By integrating the causal correlation matrix to the initial embedding, it generates a resulting feature representation that incorporates informative causal information. This enhanced feature representation can then be concatenated with the feature representations of web service traffic time series extracted by existing time series models, thereby improving prediction accuracy. Our main contributions [^3] The code, data, and other resources will be available at https://github.com/changtianluckyforever/CCMPlus. in this work can be summarized as follows:
@@
  • Method: The CCMPlus module enhances existing time series forecasting models by generating feature representations that incorporate latent causal relationships across web services, addressing a critical limitation of many previous methods. Additionally, the CCMPlus module is designed for seamless integration with existing time series forecasting models, further contributing to improved prediction accuracy.
@@
  • Experiments: Experiments conducted on three real-world web service traffic datasets (Alibaba Group, Microsoft Azure, and Ant Group) demonstrate that our method achieves superior performance in terms of Mean Squared Error (MSE) and Mean Absolute Error (MAE) compared to previous state-of-the-art methods, thereby validating the effectiveness of the CCMPlus module.
2 Related Work
2.1 Web Service Traffic Prediction
@@
Web service traffic prediction is a task critical to enabling service autoscaling, load balancing, and anomaly detection. As web service traffic is often represented as time series data, existing approaches primarily frame traffic prediction as a time series prediction problem [25, 48, 3].
@@
Early research focused on statistical prediction methods [17, 11, 18] such as Moving Average, Auto-Regression, and Autoregressive Integrated Moving Average. These methods are valued for their simplicity and interpretability but are constrained by strict stationarity requirements. Moreover, they struggle to extend to multi-dimensional or non-linear data, limiting their applicability in dynamic environments. To overcome these limitations, methods like HOPBLR [14], LLR [7] and TWRES [12] employed machine learning techniques, including Logistic Regression and Support Vector Regression, respectively, for traffic prediction. However, these approaches were hindered by the limited expressive capacity of their models, resulting in suboptimal prediction accuracy. More recent advancements have shifted towards deep learning methods. CrystalLP [29] and GRUWP [9] utilize Long Short-Term Memory networks and Gated Recurrent Units, respectively, to predict service workloads. MagicScaler [25] proposes a novel multi-scale attentive Gaussian process-based predictor, capable of accurately forecasting future demands by capturing scale-sensitive temporal dependencies. The Performer [27] integrates the Transformer architecture into an encoder-decoder paradigm for service workload prediction. It leverages the self-attention mechanism to model temporal correlations and learn both global and local representations effectively. OptScaler [48] advances this direction with a proactive prediction module comprising a long-term periodic block and a short-term local block to capture multi-scale temporal dependencies. While existing traffic prediction methods demonstrate high accuracy through advanced time series forecasting techniques, they largely neglect the underlying causal relationships within the web services. Exploring these hidden causalities holds significant potential for further improving prediction performance.
2.2 Time Series Forecasting
@@
Besides approaches specifically tailored for web service traffic prediction, general time series forecasting methods are also applied to predict web traffic [3, 48, 8].
@@
Recent advancements in neural network architectures have significantly enhanced temporal modeling capabilities. Neural network approaches for time series forecasting can be categorized into five paradigms [40, 33]: RNN-based, CNN-based, Transformer-based, MLP-based, and large language model (LLM)-based methods. Empirical methods often integrate components from the aforementioned categories, utilizing specific designs to effectively capture critical temporal features [40]. These specific designs incorporate series decomposition, multi-periodicity analysis, and multi-scale mixing architectures. For series decomposition, Autoformer [44] introduces a decomposition block based on moving averages, enabling the separation of complex temporal variations into seasonal and trend components. Building on this foundation, DLinear [45] utilizes series decomposition as a preprocessing step prior to performing linear regression. Crossformer [46] segments time series data into subseries-level patches and employs a Two-Stage Attention layer to effectively model cross-time and cross-variable dependencies within each patch. iTransformer [22] leverages the global representation of entire series and applies attention mechanisms to these series-wise representations, facilitating the capture of multivariate correlations. TimeXer [42] integrates external information into the Transformer architecture through a carefully designed embedding strategy, allowing the inclusion of external information into patch-wise representations of endogenous series. In the context of multi-periodicity, NBEATS [24] employs multiple trigonometric basis functions to model time series, providing a robust framework for handling periodic patterns. Similarly, TimesNet [43] applies Fourier Transform to decompose time series into components of varying periodic lengths and utilizes a modular architecture to process these decomposed components effectively. With respect to multi-scale mixing architectures, Pathformer [6] adopts multi-scale patch representations and applies dual attention mechanisms across these patches to capture both global correlations and local details, thereby addressing temporal dependencies comprehensively. TimeMixer [40] captures temporal features by introducing a novel multi-scale mixing architecture, which comprises two key components: Past-Decomposable-Mixing, designed to leverage disentangled series for multi-scale representation learning, and Future-Multipredictor-Mixing, which ensembles complementary forecasting skills across multi-scale series to enhance prediction accuracy.
@@
While general time series forecasting methods have been applied to web service traffic prediction [25, 48, 5], these methods often overlook causal relationships between services. In contrast, our CCMPlus module computes a causal correlation matrix used to generate temporal features incorporating causal relationships, significantly improving web service traffic prediction accuracy.
3 Background
@@
In this section, we begin by giving a formal definition of web service traffic prediction task. Subsequently, we provide an overview of the causality theory, Convergent Cross Mapping (CCM), which serves as the theoretical foundation for our proposed CCMPlus module.
3.1 Web Service Traffic
Figure 2: Web service traffic time series from Alibaba Cloud. Each point indicates the number of service requests at each time point.
@@ 3.1 Web Service Traffic
Figure 2 presents an illustrative example of web service traffic time series from Alibaba Cloud. It exhibits significant fluctuations and frequent changes, primarily driven by human behavior and activity patterns. Accurately predicting web service traffic remains a well-recognized challenge due to the inherent complexity of traffic patterns, as highlighted by the research community [25, 48]. The web traffic prediction task can be formulated as follows:
$y(t)=P(y(t-\alpha),y(t-2\alpha),\dots,y(t-k\alpha)),$ (1)
where $y(t)$ is the number of request at time $t$, $\alpha$ denotes the prediction granularity, and $k$ is the historical sequence length.
3.2 Convergent Cross Mapping
@@
Convergent Cross Mapping (CCM) theory published in Science [32] was originally proposed in the field of ecology and is designed to detect causal relationships between species. The original work includes around 50 pages of supplementary material explaining the CCM theory. The exposition of CCM is often illustrated using the context of the Lorenz system. Its trajectory forms a manifold $M$ in the state space, which consists of a collection of points that represent all possible states over time, with these points connected to create a structured geometric space.
@@
The shadow manifolds, $M_{x}$ or $M_{y}$, represents the projection of the original manifold $M$ onto the system variables $X$ or $Y$, respectively. Specifically, a lagged coordinate embedding utilizes $E$ time-lagged values of $X(t)$ as coordinate axes to reconstruct the shadow manifold $M_{x}$. A point on $M_{x}$, denoted as $\underline{x}(t)$, is an $E$-dimensional vector expressed as:
@@
where $\tau$ is a positive time lag, and $E$ denotes the embedding dimension. Similarly, the same approach can be applied to points $\underline{y}(t)$ in the shadow manifold $M_{y}$.
@@
Cross mapping refers to the process of identifying contemporaneous points in the manifold $M_{x}$ of one variable $X$ based on points in the manifold $M_{y}$ of another variable $Y$. Specifically, given a point $\underline{y}(t)$ in the manifold $M_{y}$, the corresponding point in time from the manifold $M_{x}$ is $\underline{x}(t)$. As illustrated in Figure 3, if $Y$ exerts a causal effect on $X$, information from $Y$ will be stored in $X$. Consequently, the neighbors of $\underline{x}(t)$ in $M_{x}$ will correspond to points with the same time indices in $M_{y}$, and these corresponding points will also be neighbors of $\underline{y}(t)$. However, if $Y$ has no causal effect on $X$, the information about $Y$ in $X$ will be incomplete [32]. As a result, the timely corresponding points in $M_{y}$ will diverge and no longer be neighbors of $\underline{y}(t)$.
Figure 3: The point $\underline{y}(t)$ in the manifold $M_{y}$ corresponds to the contemporaneous point in time $\underline{x}(t)$ in the manifold $M_{x}$.
@@
Convergence in CCM implies that if the variable $Y$ has a causal effect on $X$, extending the observation period improves the ability to predict $Y$ using points on the shadow manifold $M_{x}$. A longer observation period provides more trajectories to fill the gaps in the manifold, resulting in a more defined structure, which enhances the prediction of $\hat{Y}(t)\mid M_{x}$. Conversely, if two variables do not have a causal relationship, refining their manifolds will not lead to an improvement in predictive accuracy. More details of the CCM algorithm can be found in Appendix Section A.
4 Methodology